# Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization

### v1

@@ 4.2.1 Asymptotics with increasing sample size
Similar to the previous Subsection 4.1.1, we investigate the effect of varying sample size $n$ and data dimension $p$ on volatility predictions on the efficient frontier. Again we set $c=1/2$ or $c=3/2$, with $n=500,1000,1500,2000$, ${\bm{\Sigma}}={\bm{\Sigma}}_{3}$, and ${\bm{\mu}}={\bm{\mu}}_{3}$ or ${\bm{\mu}}_{4}$. Following the same procedure as described above in Section 4.2, for each $(n,p)$ and each given $\mu_{0}$, we obtain $100$ values of $\sigma_{0}^{2}$ and $\widehat{\sigma}^{2}$, denoted as $\sigma_{0,b}^{2}$ and $\widehat{\sigma}^{2}_{b}$ for $b\in[100]$. The outcomes are subsequently illustrated in Figure 4.
Figure 4: Simulation results with increasing $n$. $x$-axis in all figures shows different values of $\mu_{0}$. In Figures 4(a)-4(d), $y$-axis shows $\sum_{b=1}^{100}(\widehat{\sigma}_{b}^{2}/\sigma_{0,b}^{2}-1)^{2}/100$. Similarly, in Figures 4(e)-4(h), $y$-axis shows $\sum_{b=1}^{100}(\widehat{\sigma}_{b}^{2}-\sigma_{0,b}^{2})^{2}/100$.
The MSE’s of both the difference, $\widehat{\sigma}^{2}-\sigma_{0}^{2}$, and the ratio, $\widehat{\sigma}^{2}/\sigma_{0}^{2}-1$, are illustrated in Figure 4 as we change the sample size $n$. Note that $\mathcal{A}_{rr}$ is bounded in this simulation. Therefore, the MSE gradually decreases for each $\mu_{0}$ as $n$ grows. Figures 4(c) and 4(g) exhibit some differences at $n=1500$ and $n=2000$, but the MSEs at these sample sizes are quite similar and considerably smaller than those at $n=500$ and $n=1000$. Another interesting observation is that as $\mu_{0}$ increases, the volatility absolute error also increases, while the volatility relative error stays roughly the same. Note that Theorem 3.2 states as long as $\mu_{0}$ is within a reasonable range, the absolute and relative error converges to zero. However, when $\mu_{0}$ is large, the volatility becomes large too in the efficient frontier, and the absolute difference between $\widehat{\sigma}^{2}$ and $\sigma^{2}$ may also be large, leading to the U-shape in Figures 4(e)-4(h).
@@ 5 Real Data Analysis
In this section, we conduct real data experiments to demonstrate the effectiveness and applicability of our methodology across various scenarios. The mean-variance portfolio in Section 5.1 and the efficient frontier in Section 5.2 are both constructed using daily stock returns from the constituents of S&P500. Since estimating the mean vector is difficult, we also consider the global minimum variance portfolio in Appendix B.1, which has the advantage of no requirement about the knowledge of ${\bm{\mu}}$. In addition, we consider calibrated models in Appendix B.2 to show the effect of estimating ${\bm{\mu}}$ in the mean-variance portfolio in a more straightforward way.
After deleting stocks with missing values, we have $p=365$ stocks in total. To robustly control extreme returns, we apply a truncation at 99% quantile. Portfolios are built using historical data spanning different time length, including one year, two years, and four years, and rebalanced monthly to reflect changing market dynamics. To illustrate the portfolio construction, we give an example of building up the portfolio to be held for January 2017 using one-year data. We utilize historical data from January 2016 to December 2016 to determine the optimized portfolio positions (i.e. the allocation vector $\mathbf{w}^{*}$). These positions are then held for the entire January 2017. Then we can calculate the excess return of the portfolio $\mathbf{w}^{*}$ for each trading day in January 2017. We refer to January 2017 as the current testing month. We repeat the procedure in a rolling fashion for all testing months spanning from January 2013 to June 2023 and record the daily returns for each trading day.
@@ 5.1 Mean variance portfolio
In this section, we investigate mean-variance portfolio construction in order to achieve high Sharpe ratio. Recall that mean-variance portfolio allocation vector is $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}$. We set its scaling by assuming $\sum_{i}|\mathbf{w}_{i}|=1$, meaning that the book of the portfolio is one unit. The weight vector $\mathbf{w}$ is thus expressed as:
$\mathbf{w}=\frac{(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\|(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}\|_{1}}.$ (5.1)
In this experiment, we set $\mathbf{Q}=q\mathbf{I}$ and vary the value of $q$ from the candidate sets $\mathcal{Q}$. Based on Theorem 2.7, for each $q\in\mathcal{Q}$, we can utilize $\widehat{SR}$ to optimize the value of $q$ that yields a high $SR$. We present the steps below for our mean-variance portfolio analysis.
@@
    Consider two candidate sets $q\in\mathcal{Q}_{1}$ and $q\in\mathcal{Q}_{2}$: $\mathcal{Q}_{1}=[1:30]/5$ for $c>1$ and $\mathcal{Q}_{1}=[1:30]/1.5$ for $c<1$; $\mathcal{Q}_{2}=[1:30]/10$ considers a narrower range of values, resulting in smaller gaps thus more refined differences among values of $q$.
@@
    For each testing month, we observe the daily returns $\mathbf{R}\in\mathbb{R}^{n\times p}$, where $n$ is the total number of trading days with one-, two-, and four-year historical data length and $p=365$ is the number of selected stocks. Calculate the sample covariance matrix $\widehat{\bm{\Sigma}}$ and the portfolio weight $\mathbf{w}$ is given by (5.1).
@@
    For each testing month, we run experiments for all candidate $q$ values and also consider no regularization, i.e. $q=0$, where we have $\mathbf{w}\propto\widehat{\bm{\Sigma}}^{+}{\bm{\mu}}$ and $\widehat{\bm{\Sigma}}^{+}$ is the pseudo inverse, and the optimized $q^{*}\in\mathcal{Q}$ using the estimation in Theorem 2.7. Here,
    $q^{*}=\operatorname{argmax}_{q\in\mathcal{Q}}\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}\bigg{)}\cdot\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}{\bm{\mu}}}}.$
@@
    We roll the procedure above for all testing months. Note that the value of $q^{*}$ changes from month to month. With the weight vector $\mathbf{w}$ using all $q\in\mathcal{Q}$, $q=0$ or $q^{*}$, we can then compute the portfolio returns for each trading day in the testing month.
  5. 5.
    We report the average Sharpe ratio of daily portfolio returns over the future three years.
Our target in the experiment is to verify the effectiveness of our methodology, instead of predicting ${\bm{\mu}}$ in this paper. So we use the mean from each testing month as ${\bm{\mu}}$ to construct portfolio.
Figure 5: SR of mean-variance portfolios. The x-axis labels the rolling period, while the y-axis represents the out-of-sample SR of the portfolio returns every three years in the future. The blue solid line, orange dash-dot line and green dash line correspond to the SR with $q=0$, $q=\min_{q\in\mathcal{Q}}q$ and $q=\max_{q\in\mathcal{Q}}q$ respectively. The boxplot displays the Sharpe ratio for all $q\in\mathcal{Q}$, and the red triangle indicates the SR under our optimized $q\in\mathcal{Q}$.
Figures 5(a)-5(c) present the results obtained with $q\in\mathcal{Q}_{1}$, while Figures 5(d)-5(f) present the results obtained with $q\in\mathcal{Q}_{2}$. Obviously, if we apply $q=0$ (the blue curve), we fail to attain the best possible SR, especially when sample size is low. If we always use the constant $q=\max_{q\in\mathcal{Q}}q$ (the orange curve), it is also a bad choice of regularization. In most of the time periods, $q=\min_{q\in\mathcal{Q}}q$ (the green curve) achieves a higher SR. However, more notably, the optimized $q^{*}\in\mathcal{Q}$ (the red triangles) consistently demonstrates superior performance throughout the experiments, regardless of the candidate sets and the varying lengths of historical data. From the results, we see clear advantage of actively optimized $q^{*}\in\mathcal{Q}$ for each testing month over some ad-hoc constant $q\in\mathcal{Q}$ for all testing months. The improvement indeed comes from our newly proposed out-of-sample SR estimator.
@@ 5.2 Efficient frontier
In this section, we test the application of Theorem 3.2 to efficient frontier evaluation of real data. For each specific $\mu_{0}$, we need to estimate the standard deviation of the regularized mean-variance portfolio. We consider the optimization problem presented in (3.2). Here are the steps of our real data experiment.
@@
    Set $\mathbf{Q}=q\mathbf{I}$ and let $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ the same as in Section 5.1. We vary the values of $\mu_{0}$ from $0.4$ to $6$ with an increment of $0.4$. For each $\mu_{0}$, we repeat Steps 2-5 below.
@@
    For each testing month, centralize the observed daily returns $\mathbf{R}\in\mathbb{R}^{n\times p}$, where $n$ corresponds to one-, two-, or four-year historical data and $p=365$. Then, calculate the sample covariance matrix $\widehat{\bm{\Sigma}}$ from the observed $\mathbf{R}$.
@@
    Let $\mathbf{r}$ be the average return vector in the testing month. The optimal portfolio is given by $\mathbf{w}^{*}=\mathbf{g}+\mu_{0}\mathbf{h}$, where $\mathbf{g}$ and $\mathbf{h}$ is given by (3.5), with $\mathbf{Q}$ replaced by $q\mathbf{I}$. We run experiments for all $q$ values in the candidate sets, including the case of $q=0$ and the optimized $q^{*}$ value. By Theorem 3.2, the optimized $q^{*}$ value is obtained by minimizing $\frac{(\mathbf{g}+\mu_{0}\mathbf{h})^{\top}\widehat{\bm{\Sigma}}(\mathbf{g}+\mu_{0}\mathbf{h})}{(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1})^{2}}$ over all $q\in\mathcal{Q}$. The case $q=0$ replaces $(\widehat{\bm{\Sigma}}+q\mathbf{I})^{-1}$ with $\widehat{\bm{\Sigma}}^{+}$ in $\mathbf{w}^{*}$.
@@
    We repeat Step 2 for each testing month from January 2013 to June 2023. This allows us to collect daily portfolio returns for each $q$ value.
  5. 5.
    We calculate the standard deviation of the daily returns for each $q$ value, including $q=0$ and $q=q^{*}$, over the ten-year period. This allows us to generate a boxplot of the standard deviation for each candidate $q$ value in $\mathcal{Q}_{1}$ or $\mathcal{Q}_{2}$ given $\mu_{0}$. We also obtain the standard deviation of the portfolios corresponding to $q=0$ and $q=q^{*}$.
Figure 6: The corrected efficient frontier. The x-axis represents the value of $\sigma$, while the y-axis represents different values of $\mu_{0}$. The blue solid line, orange dash-dot line and green dash line in the figure correspond to the different values of $\sigma$ with $q=0$, $q=\min_{q\in\mathcal{Q}}q$ and $q=\max_{q\in\mathcal{Q}}q$ respectively. The boxplot displays the values of $\sigma$ for all $q\in\mathcal{Q}$, and the red triangle indicates the volatility of the portfolio returns under our optimized $q\in\mathcal{Q}$.
As shown in Figure 6, the standard approach with no regularization ($q=0$) substitutes $(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}$ with $\widehat{\bm{\Sigma}}^{+}$ and exhibits a significantly larger volatility in comparison to the regularized portfolios for any given $\mu_{0}$. Furthermore, as illustrated in Figures 6(d) and 6(f), the minimum standard deviation occurs when $q$ is approximately $\max_{q\in\mathcal{Q}_{2}}q$ and $\min_{q\in\mathcal{Q}_{2}}q$ respectively. However, our corrected estimator consistently identify the optimal $q\in\mathcal{Q}_{2}$ and result in a smaller standard deviation. This again demonstrates the practical usefulness of Theorem 3.2 in real efficient frontier recovery of the regularized portfolios.

### v3

@@ 4.2.1 Asymptotics with increasing sample size
Similar to the previous Subsection 4.1.1, we investigate the effect of varying sample size $n$ and data dimension $p$ on volatility predictions on the efficient frontier. Again we set $c=1/2$ or $c=3/2$, with $n=500,1000,1500,2000$, ${\bm{\Sigma}}={\bm{\Sigma}}_{3}$, and ${\bm{\mu}}={\bm{\mu}}_{3}$ or ${\bm{\mu}}_{4}$. Following the same procedure as described above in Section 4.2, for each $(n,p)$ and given $\mu_{0}$, we obtain $1000$ values of $\sigma_{0}^{2}$ and $\widehat{\sigma}^{2}$, denoted as $\sigma_{0,b}^{2}$ and $\widehat{\sigma}^{2}_{b}$ for $b\in[1000]$. The outcomes are subsequently illustrated in Figure 4. The MSE’s of the ratio $\widehat{\sigma}^{2}/\sigma_{0}^{2}-1$, the SR difference $\mu_{0}/\widehat{\sigma}-\mu_{0}/\sigma_{0}$, and the variance difference $\widehat{\sigma}^{2}-\sigma_{0}^{2}$ are illustrated in Figure 4 as we change the sample size $n$. In Figure 4, we see that the discrepancy between $\sigma_{0}$ and $\widehat{\sigma}$ decreases as the sample size $n$ increases as expected. The ratio consistently remains around 1, as shown in Figures 4(a)-4(d), supporting the ratio consistency in Theorem 3.2. When examining Sharpe ratio differences, Figures 4(e) and 4(g) show that the differences increase as $\mu_{0}$ grows, whereas Figures 4(f) and 4(h) show small differences even for larger $\mu_{0}$. This aligns with our theoretical result in Item 1 of Theorem 3.2 as $\mathcal{A}_{rr}$ is unbounded for ${\bm{\mu}}={\bm{\mu}}_{3}$ and bounded for ${\bm{\mu}}={\bm{\mu}}_{4}$. Regarding differences between $\widehat{\sigma}^{2}$ and $\sigma_{0}^{2}$, Figures 4(i) and 4(k) show small differences for larger $\mu_{0}$ when $\mathcal{A}_{rr}$ is unbounded so that $\mu_{0}$ can take a large value and still ensures the convergence of $\widehat{\sigma}^{2}$ to $\sigma_{0}^{2}$. In contrast, Figures 4(j) and 4(l) show larger differences for larger $\mu_{0}$ when $\mathcal{A}_{rr}$ is bounded. This matches our result in Item 2 of Theorem 3.2.
Figure 4: Simulation results with increasing $n$. $x$-axis in all figures shows different values of $\mu_{0}$. In Figures 4(a)-4(d), $y$-axis shows $\sum_{b=1}^{1000}(\widehat{\sigma}_{b}^{2}/\sigma_{0,b}^{2}-1)^{2}/{1000}$. In Figures 4(e)-4(h), $y$-axis shows $\sum_{b=1}^{1000}(\frac{\mu_{0}}{\widehat{\sigma}_{b}}-\frac{\mu_{0}}{\sigma_{0,b}})^{2}/{1000}$. And in Figures 4(i)-4(l), $y$-axis shows $\sum_{b=1}^{1000}(\widehat{\sigma}_{b}^{2}-\sigma_{0,b}^{2})^{2}/{1000}$.
@@ 5 Real Data Analysis
In this section, we conduct real data experiments to demonstrate the effectiveness and applicability of our methodology across various scenarios. The MV portfolio in Section 5.1 and the efficient frontier in Section 5.2 are both constructed using daily stock returns from the constituents of S&P500 assuming we know the out-of-sample ${\bm{\mu}}$. In addition, we consider the global minimum variance portfolio in Appendix B.1, which has the advantage of not requiring the knowledge of ${\bm{\mu}}$, and calibrated models in Appendix B.2 to show the effect of estimating ${\bm{\mu}}$ in the MV portfolio. Last but not least, we consider the MV portfolio with historical sample mean $\bm{\widehat{\mu}}$ in Appendix B.3 following Remark 2.9.
After deleting stocks with missing values, we end up $p=365$ stocks in total. Portfolios are built using historical data spanning different time length, including one year, two years, and four years, and rebalanced monthly to reflect changing market dynamics. To illustrate the portfolio construction, we give an example of building up the portfolio for Jan 2017 using one-year data. We utilize historical data from Jan 2016 to Dec 2016 to determine the optimized portfolio positions (i.e. the allocation vector $\mathbf{w}^{*}$). These positions are then held for the entire Jan 2017. We refer to Jan 2017 as the current testing month and calculate the excess returns of the portfolio $\mathbf{w}^{*}$ for each trading day in Jan 2017. We repeat the procedure in a rolling fashion for all testing months spanning from Jan 2013 to Jun 2023 and record the daily returns for each trading day. The candidate set for the regularization matrix $\mathcal{Q}_{1}$ is defined as $\mathcal{Q}_{1}=\{q\cdot\widehat{\bm{\Sigma}}_{pre},q\in[1:30]/10\}$, where $\widehat{\bm{\Sigma}}_{pre}$ represents the sample covariance matrix of returns calculated from Jan 2004 to Dec 2008, which is a pre-training period not overlapping with data for portfolio construction and evaluation. We also consider a second candidate set $\mathcal{Q}_{2}=\{q\cdot\mathbf{I}_{p},q\in[1:30]/10\}$, where $\mathbf{I}_{p}$ is the identity matrix of dimension $p$.
@@ 5.1 Mean variance portfolio
In this section, we construct the MV portfolios in order to achieve high Sharpe ratio. Recall that the MV portfolio allocation vector is $\mathbf{w}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}$. We set its scaling by assuming $\sum_{i}|\mathbf{w}_{i}|=1$, meaning that the book of the portfolio is one unit. The weight vector $\mathbf{w}$ is
$\mathbf{w}={(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}/{\|(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}\|_{1}}.$ (5.1)
Based on Theorem 2.7, we can utilize $\widehat{SR}$ to optimize $\mathbf{Q}\in\mathcal{Q}_{1}$ or $\mathbf{Q}\in\mathcal{Q}_{2}$ to yield a high $SR$. We present the steps for our MV portfolio analysis.
@@
    For each testing month, we observe the historical daily returns $\mathbf{R}\in\mathbb{R}^{n\times p}$, where $n$ is the total number of trading days with one-, two-, or four-year data and $p=365$ is the number of selected stocks. Calculate $\widehat{\bm{\Sigma}}$ and the portfolio $\mathbf{w}$ by (5.1).
@@
    For each testing month, we run experiments for all candidate $\mathbf{Q}\in\mathcal{Q}$ and also consider no regularization, i.e. $\mathbf{Q}=\text{\bf 0}$, where we have $\mathbf{w}\propto\widehat{\bm{\Sigma}}^{+}{\bm{\mu}}$ and $\widehat{\bm{\Sigma}}^{+}$ is the pseudo inverse, and the optimized $\mathbf{Q}^{*}\in\mathcal{Q}$ using the estimation in Theorem 2.7. Here,
    $\mathbf{Q}^{*}=\operatorname{argmax}_{\mathbf{Q}\in\mathcal{Q}}\bigg{(}1-\frac{c}{p}{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\bigg{)}\cdot\frac{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}{\sqrt{{\bm{\mu}}^{\top}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}}}.$
@@
    We roll the procedure above for all testing months. Note that the value $\mathbf{Q}^{*}$ changes from month to month. With the weight vector $\mathbf{w}$ using all $\mathbf{Q}\in\mathcal{Q}$, $\mathbf{Q}=\text{\bf 0}$ or $\mathbf{Q}^{*}$, we can then compute the portfolio returns for each trading day in the testing month.
@@
    We report the realized Sharpe ratio of daily portfolio returns over the future three years.
Figure 5: SR of mean-variance portfolios. The x-axis labels the rolling period, while the y-axis represents the out-of-sample SR of the portfolio returns every three years in the future. The blue solid line, orange dash-dot line and green dash line correspond to the SR with $q=0$, $q=0.1$ (minimum $q$) and $q=3$ (maximum $q$) respectively. The boxplot displays SR for all $\mathbf{Q}\in\mathcal{Q}$, and the red triangle indicates the SR under our optimized $\mathbf{Q}^{*}\in\mathcal{Q}$.
In this experiment, we use the mean from each out-of-sample testing month as ${\bm{\mu}}$ to construct MV portfolio. Readers may refer to Appendix B.3 for real data experiments with unknown ${\bm{\mu}}$. Figures 5(a)-5(c) present the results obtained with $\mathbf{Q}\in\mathcal{Q}_{1}$, while Figures 5(d)-5(f) present the results obtained with $\mathbf{Q}\in\mathcal{Q}_{2}$. Obviously, when sample size is small, if we apply $\mathbf{Q}=\text{\bf 0}$ (the blue curve) or $q=0.1$ (the orange curve, minimum $q$), we fail to attain the best possible SR. If we use $q=3$ (the green curve, maximum $q$), it is also a bad choice when sample size is large. However, the optimized $\mathbf{Q}^{*}\in\mathcal{Q}$ (the red triangles) consistently demonstrates superior performance throughout the experiments, regardless of the candidate sets and the varying lengths of historical data. From the results, we see clear advantage of actively optimized $\mathbf{Q}^{*}\in\mathcal{Q}$ for each testing month over some ad-hoc constant $\mathbf{Q}\in\mathcal{Q}$ for all testing months. Moreover, the optimal regularization $\mathbf{Q}^{*}\in\mathcal{Q}_{1}$ performs better than $\mathbf{Q}^{*}\in\mathcal{Q}_{2}$. For instance, comparing Figures 5(c) and 5(f), using the identical blue lines as the benchmark, the optimal SR values (red triangles) obtained from $\mathcal{Q}_{1}$ are visibly higher than those obtained from $\mathcal{Q}_{2}$.
@@ 5.2 Efficient frontier
In this section, we test the real data application of Theorem 3.2 to efficient frontier evaluation. For each specific $\mu_{0}$, we need to estimate the standard deviation of the regularized MV portfolio, presented in (3.2). Here are the steps of our real data analysis.
@@
    For each testing month, centralize the observed historicla daily returns $\mathbf{R}\in\mathbb{R}^{n\times p}$, where $n$ corresponds to one-, two-, or four-year data and $p=365$. Then, calculate the sample covariance $\widehat{\bm{\Sigma}}$ from the observed $\mathbf{R}$. Let $\mathbf{r}$ be the average return vector in the testing month. The optimal portfolio is given by $\mathbf{w}^{*}=\mathbf{g}+\mu_{0}\mathbf{h}$, where $\mathbf{g}$ and $\mathbf{h}$ is given by (3.5).
@@
    We run experiments for all $\mathbf{Q}$ in the candidate sets, including $\mathbf{Q}=\text{\bf 0}$ and the optimized $\mathbf{Q}^{*}$. By Theorem 3.2, the optimized $\mathbf{Q}^{*}$ is obtained by minimizing $\frac{(\mathbf{g}+\mu_{0}\mathbf{h})^{\top}\widehat{\bm{\Sigma}}(\mathbf{g}+\mu_{0}\mathbf{h})}{(1-c/p\cdot{\mathrm{tr}}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1})^{2}}$ over all $\mathbf{Q}\in\mathcal{Q}$. The case $\mathbf{Q}=\text{\bf 0}$ replaces $(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}$ with $\widehat{\bm{\Sigma}}^{+}$ in $\mathbf{w}^{*}$.
@@
    We repeat Step 2 for each testing month from Jan 2013 to Jun 2023. This allows us to collect daily portfolio returns for each $\mathbf{Q}$.
@@
    We calculate the standard deviation of the daily returns for each $\mathbf{Q}$, including $\mathbf{Q}=\text{\bf 0}$ and $\mathbf{Q}^{*}$, over the ten-year period. This allows us to generate a boxplot of the standard deviation for each candidate $\mathbf{Q}$ in $\mathcal{Q}_{1}$ or $\mathcal{Q}_{2}$ given $\mu_{0}$.
Figure 6: The corrected efficient frontier. The x-axis represents the values of $\sigma$, while the y-axis represents values of $\mu_{0}$. The blue solid line, orange dash-dot line and green dash line in the figure correspond to $q=0$, $q=0.1$ (minimum $q$) and $q=3$ (maximum $q$) respectively. The boxplot displays the values of $\sigma$ for all $\mathbf{Q}\in\mathcal{Q}$, and the red triangle indicates the volatility of the portfolio returns under our optimized $\mathbf{Q}^{*}\in\mathcal{Q}$.
As shown in Figure 6, the standard approach with no regularization ($\mathbf{Q}=\text{\bf 0}$) substitutes $(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}$ with $\widehat{\bm{\Sigma}}^{+}$ and exhibits a larger volatility in comparison to the regularized portfolios for any given $\mu_{0}$. Furthermore, as illustrated in Figures 6(d) and 6(f), the minimum standard deviation occurs when $q$ is approximately $3$ and $0.1$ respectively. However, our corrected estimator consistently identify the optimal $\mathbf{Q}\in\mathcal{Q}_{2}$ and result in a smaller standard deviation. This again demonstrates the practical usefulness of Theorem 3.2 in real efficient frontier recovery of the regularized portfolios.